when invested at an annual interest rate of 7.5 an account earned $217.50 of simple interest in one year. how much money was originally invested in the account?

Answer by stanbon(75887)

when invested at an annual interest rate of 7.5 an account earned $217.50 of simple interest in one year. how much money was originally invested in the account?
------
I = P*r*t
217.5 = P*0.075
----
P = 217.5/0.075
P = $2900
---------------
Cheers,
Stan H.
------------